Solve: 4/7x-3/2=x-1/7
Answer by stanbon(75887)

4/7x-3/2=x-1/7
------
Multiply thru by 14 to get:
8x - 21 = 14x - 2
-----
6x = -19
-----
x = -19/6
=============
Cheers,
Stan H.
=============